# FaceXFormer: A Unified Transformer for Facial Analysis

### v2

@@ 4.1 Datasets and Metrics
We perform co-training, where the model is simultaneously trained for multiple tasks using a total of 9 datasets with task-specific annotations. We conduct intra-dataset and cross-dataset evaluations and present our results on the test sets according to the standard protocol for each task using the following datasets:
Training: Face Farsing: CelebAMaskHQ [38]; Landmarks Detection: 300W [75]; Head Pose Estimation: 300W-LP [128]; Attributes Prediction: CelebA [50]; Facial Expression Recognition: RAF-DB [42], AffectNet [58]; Age/Gender/Race estimation: UTKFace [120], FairFace [32]; Visibility Prediction: COFW [7].
Test (Intra-dataset): Face Parsing: CelebAMaskHQ [38]; Landmarks Detection: 300W [128]; Attributes Prediction: CelebA [50]; Facial Expression Recognition: RAF-DB [42]; Age/Gender/Race Estimation: UTKFace [120], FairFace [32]; Visibility Prediction: COFW [7].
Test (Cross-dataset): Landmarks Detection: 300VW [78]; Head Pose Estimation: BIWI [18]; Attributes Prediction: LFWA [101].
The evaluation metrics used are the F1-score for face parsing, Normalized Mean Error (NME) for landmark prediction, Mean Absolute Error (MAE) for head pose estimation and age estimation, accuracy for facial expression recognition, attributes prediction, gender estimation, race estimation, and recall at $80$% precision for visibility prediction.
@@
Table 3: Performance comparison on facial expression recognition, face visibility prediction and age estimation. $\times$ indicates a model that doesn’t perform the task.
@@ 4.2 Implementation Details
We train our models using a distributed PyTorch setup on eight A6000 GPUs, each equipped with $48$GB of memory. The models’ backbones are initialized with ImageNet pre-trained weights and processes input images at a resolution of $224\times 224$. We employ the AdamW optimizer with a weight decay of $1e^{-5}$. All models are trained for $12$ epochs with a batch size of $48$ on each GPU, and an initial learning rate of $1e^{-4}$, which decays by a factor of $10$ at the $6^{th}$ and $10^{th}$ epochs. We train the model for three additional epochs for some tasks. For data augmentation, we randomly apply Gaussian blur, grayscale conversion, gamma correction, occlusion, horizontal flipping, and affine transformations, such as rotation, translation and scaling. The number of FaceX decoder $N$ is set to two. To ensure stable training across tasks when using multiple datasets of varying sample sizes, we equalize the representation of each task’s samples in every batch through upsampling. Additional details on our implementation are provided in the supplementary.
@@ 4.3 Main results
In Tab. 2, Tab. 4, Tab. 3, we present a comparative analysis of FaceXFormer with recent methods across a variety of tasks. A significant highlight of our work is its unique capability to deliver promising results across multiple tasks using a single unified model. Specifically, FaceXFormer achieves state-of-the-art performance in face parsing, with a mean F1 score of $92.01$ on CelebAMaskHQ at a resolution of $224\times 224$, which is half the input size required by other state-of-the-art methods. Furthermore, it demonstrates superior performance in head pose estimation and landmark detection, achieving a mean MAE of $3.52$ and a mean NME of $4.67$, respectively. Additionally, FaceXFormer provides a significant performance boost in attributes prediction and visibility prediction, achieving an accuracy of $91.83$% on the CelebA dataset and $72.56$% on COFW. It also performs competitively in age estimation, achieving the second-best score of $4.17$, and achieves an accuracy of $88.24$% in facial expression recognition. The results on gender estimation across different race categories is shown in Tab. 6. We present the addtional cross-dataset results in Appendix B.
Recent models such as SwinFace [66], QFace [83], and Faceptor [67] also aim to unify tasks but only address a subset of them. These models tend to exclude complex tasks such as segmentation, head pose estimation, and landmark prediction due to their conflicting training objectives. In contrast, FaceXFormer, with its learnable task-specific tokens, seamlessly unifies these tasks and achieves state-of-the-art performance across them. It outperforms previous multi-task models in tasks such as segmentation, head pose estimation, landmark prediction, attributes prediction, and visibility prediction, while achieving the second-best performance in age estimation. In this work, we simultaneously train for nine heterogeneous tasks, presenting a more formidable challenge than previous approaches. This difficulty arises primarily from the distinct and diverse nature of the tasks, which require task-specific features, pre-processing steps, and often conflicting training objectives. Despite these challenges, FaceXFormer effectively handles multiple tasks, delivering SOTA or competitive performance, thereby establishing itself as a unified SOTA model.
@@ 4.4 Qualitative “in-the-wild” results
In this section, we present the qualitative results of FaceXFormer on randomly selected “in-the-wild” images. We select three random images and showcase the results for face parsing, head pose estimation, landmarks prediction, age estimation, gender and race classification, and attributes prediction in Figure 3. Notably, the model successfully performs complex tasks such as face segmentation, head pose estimation, and landmark prediction, even when the input samples are out-of-distribution. Furthermore, FaceXFormer can be effectively used as a tool to generate multiple annotations for each image, making it valuable for various downstream tasks. These results highlight FaceXFormer’s robust performance in challenging, real-world scenarios.
@@ 5 Ablation Study
Table 5: Effect of different backbones on performance and FPS.
In this section, we explore the impact of various backbones and their sizes on performance. Additionally, we highlight that the proposed model exhibits minimal bias compared to other models by performing age and gender prediction across different various. We provide additional ablation of different components in our model in the supplementary.
5.1 Effect of different backbone and different size.
@@
In Table 5, we present experiments analyzing the impact of different backbones and their sizes on FaceXFormer. We group head pose estimation, landmarks prediction, and age estimation into the regression (Reg) category, while attributes prediction and facial expression recognition are categorized as classification (Cls). ConvNeXt achieves the best performance in segmentation with an F1 score of $92.08$%. The Swin Transformer backbone excels in both regression and classification tasks, with a mean error of $4.12$ and a mean accuracy of $90.03$%, respectively. In contrast, MobileNet demonstrates the lowest performance metrics, including an F1 score of $91.21$% and a mean error of $4.64$, highlighting its limitations in handling larger, more complex datasets due to its smaller receptive field compared to the Swin Transformer. The selection of the Swin Transformer as the backbone for FaceXFormer is driven by its superior scalability and global contextual understanding, both of which are essential for facial analysis tasks. Furthermore, FaceXFormer achieves real-time performance at $33.21$ FPS.
Figure 3: FaceXFormer predictions on “in-the-wild” images
5.2 Bias Analysis and Ethical Considerations
In our work, we utilize a total of $12$ datasets for training and evaluation. We obtained these datasets following the procedures stated on their respective pages and signed the license agreements if and when necessary. As we train our models on multiple datasets designed for different tasks, the subjects across different age groups, genders, and races is not equal. This imbalance may introduce bias in the model. Therefore, we provide an analysis using the FairFace [31] dataset, which is balanced in terms of age, gender and race. We follow [68] and define the ”Non-white” group to include multiple racial categories: ”Black”, ”Indian”, ”East Asian”, ”Southeast Asian”, ”Middle Eastern” and ”Latino”. As can be seen from Table 6, FaceXFormer shows the smallest performance discrepancy across different racial groups and exhibits minimal bias compared to other models despite being trained on fewer data points. This can be attributed to race estimation being the task in co-training.
Model Data Points White Non-white Average Discrepancy
Age FairFace 100K 60.05 60.63 60.52 0.58
CLIP 400M 62.25 61.95 62.00 -0.30
FaRL 20M 61.49 61.84 61.78 0.35
FaceXFormer 400K 58.94 59.44 59.34 0.50
Gender FairFace 100K 94.15 94.41 94.36 0.26
CLIP 400M 94.87 95.78 95.61 0.91
FaRL 20M 95.16 95.77 95.65 0.61
FaceXFormer 400K 95.34 95.19 95.22 -0.09
Table 6: Age and gender accuracy w.r.t race groups on FairFace
6 Discussions
Broader Impact: The world is moving towards transformers because of its potential to model large amounts of data [6, 34, 5]. Presently, the face community lacks large-scale annotated datasets to train foundational models capable of performing a wide spectrum of facial tasks. The largest clean dataset, WebFace42M [130], lacks annotations for face parsing, landmarks detection and attributes recognition. FaceXFormer can be used as an annotator for large-scale data, and can be continually improved through successive rounds of annotation and fine-tuning. We aim to propel the face community towards developing foundation models that cater to a variety of facial tasks. Additionally, FaceXFormer is a lightweight model that provides real-time output based on task-specific queries and can be appended with existing facial systems to provide additional information. It can also serve as a valuable tool in surveillance, and provide auxiliary information for subject analysis and image retrieval.
Limitations: We recognize certain limitations of the proposed FaceXFormer, particularly the requirement for complete retraining to add a new task, which may reduce its flexibility. Furthermore, although it includes token support for multiple tasks, it lacks interactivity and full promptability. Nonetheless, FaceXFormer is distinct in its ability to handle up to $9$ heterogeneous tasks within a single model, achieving state-of-the-art or competitive performance across these tasks in real-time, making it well-suited for deployment. Future work will focus on developing a large-scale pre-trained foundational model with zero-shot capabilities.
7 Conclusion
In conclusion, the FaceXformer introduces a novel end-to-end unified model that efficiently handles a comprehensive range of facial analysis tasks in real-time. By adopting a transformer-based encoder-decoder architecture and treating each task as a learnable token, our approach successfully integrates multiple tasks within a single framework. The proposed parameter-efficient decoder, FaceX, enhances the model’s ability to learn robust and generalized face representations across diverse tasks. Our comprehensive experiments demonstrate that the proposed model achieves state-of-the-art performance across multiple facial analysis tasks. Additionally, training on multiple datasets leads to better representation learning. In conclusion, we demonstrate that facial tasks can be treated as tokens, leading to the unification of tasks; following this, we hope our work provides a foundation for developing large models capable of performing multiple facial analysis tasks.
8 Acknowledgment
Appendix
Appendix A Overview
  • Cross-dataset Evaluation (Section B)
  • Ablation study (Section C)
  • In-the-wild Visualization (Section D)
  • Dataset details (Section E)
Appendix B Cross-Dataset Evaluation
@@
We conduct additional cross-dataset experiments to demonstrate the effectiveness of FaceXFormer in scenarios that closely resemble real-life conditions. These scenarios involve previously unseen, unconstrained face images characterized by significant variability in background, lighting, pose, and other factors. As shown in Table B.1, FaceXFormer outperforms the existing state-of-the-art model, STARLoss [125], on the 300VW dataset. This highlights FaceXFormer’s effectiveness in landmark detection under in-the-wild scenarios. The cross-dataset results support the rationale presented in this paper: the necessity of a unified facial analysis model capable of performing multiple tasks on unconstrained, in-the-wild faces, particularly for real-time applications. FaceXFormer addresses this gap and achieves state-of-the-art performance.
@@
Table B.1: Cross Dataset evaluation of FaceXFormer.
Appendix C Ablation Study
To evaluate the contribution of each component in FaceXFormer, we conduct an ablation study focusing on the importance of specific design choices and their impact on performance across various tasks. Specifically, we perform experiments by: a) Excluding multi-scale features and relying solely on the final-layer features. b) Removing the FXDec decoder from the architecture. The results of these experiments are summarized in Table C.1.
Method HPE Lnd Attr. Age
F1 MAE Acc. MAE
w/o multi-scale 3.70 4.70 91.21 4.21
w/o FXDec 17.16 31.49 79.90 16.15
FaceXFormer 3.52 4.67 91.83 4.17
Table C.1: Impact of various components of FaceXFormer on performance.
From the results, it is evident that the FXDec decoder, which incorporates self-attention and interactions between face tokens and task tokens, plays a critical role in performance. Without FXDec, there is a significant drop across all tasks. For instance, the Mean Absolute Error (MAE) for landmark detection increases from $4.67$ to $31.49$, and the accuracy for attribute classification drops from $91.83$% to $79.90$%. This substantial decline highlights the decoder’s importance in effectively capturing complex feature relationships necessary for these tasks. Similarly, excluding multi-scale features leads to reduced performance across all tasks, with a particularly notable impact on head pose estimation and age estimation. The F1 score for head pose estimation increases from $3.52$ to $3.70$ (indicating worse performance since lower is better), and the MAE for age estimation increases from $4.17$ to $4.21$. These results shows the importance of integrating multi-scale features to capture both global and local information essential for accurate predictions.
Appendix D In-the-wild Visualization
We randomly selected images from the publicly available CelebAMask-HQ [38] dataset and treated them as “in-the-wild” images for tasks it was not specifically trained on. We present the qualitative results for all the tokens in Figure D.1. Our observations indicate that FaceXFormer produces promising results. However, we noted inconsistencies in age estimation and race prediction.
Figure D.1: Visualization of ”in-the-wild” images for multiple tasks. Attributes represent the $40$ binary attributes defined in the CelebA [50] dataset, indicating the presence ($1$) or absence (00) of specific facial attributes.
Appendix E Datasets and Implementation Details

### v3

@@ 4.1 Datasets and Metrics
We perform co-training, where the model is simultaneously trained for multiple tasks using a total of 10 datasets with task-specific annotations. We conduct a comprehensive evaluation, comparing our approach with both task-specific and multi-task models. We present our results on the test sets according to the standard protocol for each task using the following datasets:
Train: Face Farsing: CelebAMaskHQ [44]; Landmarks Detection: 300W [82]; Head Pose Estimation: 300W-LP [138]; Attributes Prediction: CelebA [56]; Facial Expression Recognition: RAF-DB [47], AffectNet [64]; Age/Gender/Race estimation: UTKFace [128], FairFace [37]; Face Recognition: MS1MV3 [26]; Visibility Prediction: COFW [8].
Test: Face Parsing: CelebAMaskHQ [44]; Landmarks Detection: 300W [138], 300VW [86]; Head Pose Estimation: BIWI [21]; Attributes Prediction: CelebA [56], LFWA [110]; Facial Expression Recognition: RAF-DB [47]; Age/Gender/Race Estimation: UTKFace [128], FairFace [37]; Face Recognition: LFW [32], CFP-FP [84], AgeDB [65], CALFW [132], CPLFW [131] ; Visibility Prediction: COFW [8].
The evaluation metrics used are the F1-score for face parsing, Normalized Mean Error (NME) for landmark prediction, Mean Absolute Error (MAE) for head pose estimation and age estimation, accuracy for facial expression recognition, attributes prediction, gender estimation, race estimation, 1:1 verification accuracy for face recognition, and recall at $80$% precision for face visibility prediction.
@@
Table 4: Performance comparison on facial expression recognition, face visibility prediction and age estimation.
@@ 4.2 Implementation Details
We train our models using a distributed PyTorch setup on eight A6000 GPUs, each equipped with $48$GB of memory. The models’ backbones are initialized with ImageNet pre-trained weights and processes input images at a resolution of $224\times 224$. We employ the AdamW optimizer with a weight decay of $1e^{-5}$. All models are trained for $12$ epochs with a batch size of $48$ on each GPU, and an initial learning rate of $1e^{-4}$, which decays by a factor of $10$ at the $6^{th}$ and $10^{th}$ epochs. We train the model for three additional epochs for some tasks. For data augmentation, we randomly apply Gaussian blur, grayscale conversion, gamma correction, occlusion, horizontal flipping, and affine transformations, such as rotation, translation and scaling. The number of FaceX decoder $N$ is set to two. To ensure stable training across tasks when using multiple datasets of varying sample sizes, we equalize the representation of each task’s samples in every batch through upsampling. Additional details on our implementation are provided in the Appendix F.
@@ 4.3 Main results
In Table 3, Table 4, Table 6, Table 5, we present a comparative analysis of FaceXFormer against recent methods across a variety of tasks. A key highlight of our work is its unique capability to deliver promising results across multiple tasks at real-time inference speed using a single unified model. Specifically, FaceXFormer achieves state-of-the-art performance in face parsing, with a mean F1 score of $92.01$ on CelebAMaskHQ at a resolution of $224\times 224$, which is half the input size required by other state-of-the-art methods. Furthermore, it demonstrates superior performance in head pose estimation and landmark detection, achieving a mean MAE of $3.52$ and a mean NME of $4.67$, respectively. Additionally, FaceXFormer provides a significant performance boost in attributes prediction and visibility prediction, achieving an accuracy of $91.83$% on the CelebA dataset and $72.56$% on COFW. It also performs competitively in age estimation, achieving the second-best score of $4.17$, and achieves an accuracy of $88.24$% in facial expression recognition. In face recognition, FaceXFormer outperforms Faceptor, achieving a mean accuracy of $95.94$% compared to $95.28$%. However, we observe that multi-task models generally underperform compared to specialized ones in this task. This can be attributed to conflicting training objectives, which force the model to learn identity-invariant features rather than identity-specific representations crucial for accurate recognition. The results on gender estimation across different race categories is shown in Table 8. We present additional cross-dataset results in Appendix D.
Method LFW CFP-FP AgeDB CALFW CPLFW Mean
CosFace [100] 99.81 98.12 98.11 95.76 92.28 96.81
ArcFace [16] 99.83 98.27 98.28 95.45 92.08 96.78
VPL-ArcFace [17] 99.83 99.11 98.60 96.12 93.45 97.42
AdaFace [39] 99.83 99.11 98.17 96.02 93.93 97.41
SwinFace [73] 99.87 98.60 98.15 96.10 93.42 97.22
QFace [91] $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Faceptor [74] 99.40 96.34 93.65 94.75 92.27 95.28
FaceXFormer 99.68 96.75 96.35 95.50 91.41 95.94
Table 6: Performance comparison for face recognition.
Recent models such as SwinFace [73], QFace [91], and Faceptor [74] also aim to unify multiple tasks but only address a subset of them. These approaches often exclude complex tasks such as segmentation, head pose estimation, and landmark prediction. Moreover, they rely on multiple decoders and computationally expensive attention mechanisms, adding to the overall computational overhead. In contrast, FaceXFormer seamlessly unifies these complex tasks using a lightweight decoder and achieves state-of-the-art performance across them at a real-time FPS of $33.21$. It outperforms previous multi-task models in segmentation, head pose estimation, landmark prediction, attribute prediction, and face visibility prediction, while achieving the second-best performance in age estimation. In this work, we simultaneously train for ten heterogeneous tasks, posing a more formidable challenge than previous approaches. Despite this, FaceXFormer effectively handles multiple tasks, achieving SOTA or competitive performance in real time. This success can be attributed to the efficiency of the proposed lightweight decoder, which employs a novel bi-directional cross-attention mechanism.
@@ 4.4 Qualitative “in-the-wild” results
In this section, we present the qualitative results of FaceXFormer on randomly selected “in-the-wild” images. We select four random images and showcase the results for face parsing, head pose estimation, landmarks prediction, age estimation, gender and race classification, and attributes prediction in Figure 3. Notably, the model successfully performs complex tasks such as face segmentation, head pose estimation, and landmark prediction, even when input samples exhibit extreme poses, occlusions, or blurring. Furthermore, FaceXFormer can be effectively used as a tool to generate multiple annotations for each image, making it valuable for various downstream tasks. These results highlight FaceXFormer’s robust performance in challenging, real-world scenarios.
@@ 5 Ablation Study
MLP Self Cross Bi-dir. HPE Lnd Attr. Age
Fusion Attn. Attn. CA (MAE) (NME) (Acc.) (MAE)
3.70 4.70 91.21 4.21
17.16 31.49 79.90 16.15
4.63 5.30 89.98 4.84
3.52 4.67 91.83 4.17
Table 7: Impact of various components on performance.
In this section, we explore the impact of different components of FaceXFormer on performance. Additionally, we demonstrate that the proposed model exhibits minimal bias compared to other models by evaluating age and gender prediction across various demographics. Furthermore, we analyze the computational performance of different components of FaceXFormer and compare it with existing multi-task models. Additional ablation studies on the impact of using different backbones of varying sizes are provided in the Appendix C.
5.1 Impact of various components in FaceXFormer
To evaluate the contribution of each component in FaceXFormer, we conduct an ablation study focusing on the importance of specific design choices and their impact on performance across various tasks. The results of these experiments are summarized in Table 7. We observe that without MLP fusion (row 1), there is a drop in performance, highlighting the importance of of integrating multi-scale features to capture both global and local information essential for accurate predictions. The model performs extremely poorly (row 2) without cross-attention in the decoder, which is expected, as there is no interaction between face tokens and task tokens in this case. Introducing the proposed bi-directional cross-attention (row 4), which corresponds to FaceXFormer, in the decoder provides a significant boost compared to using standard cross-attention (row 3), yielding improvements of $1.11$ MAE in head pose estimation, $0.63$ NME in landmark detection, $1.85$ accuracy points in attribute prediction, and $0.67$ MAE in age estimation. These results demonstrate the importance of MLP fusion and bi-directional cross-attention in the FaceXFormer architecture.
Figure 3: FaceXFormer predictions on “in-the-wild” images
5.2 Bias Analysis and Ethical Considerations
In our work, we utilize $17$ unique datasets for training and evaluation. We obtained these datasets following the procedures stated on their respective pages and signed the license agreements if and when necessary. As we train our models on multiple datasets designed for different tasks, the subjects across different age groups, genders, and races is not equal. This imbalance may introduce bias in the model. Therefore, we provide an analysis using the FairFace [36] dataset, which is balanced in terms of age, gender and race. We follow [75] and define the ”Non-white” group to include multiple racial categories: ”Black”, ”Indian”, ”East Asian”, ”Southeast Asian”, ”Middle Eastern” and ”Latino”. As can be seen from Table 8, FaceXFormer shows the smallest performance discrepancy across different racial groups and exhibits minimal bias compared to other models despite being trained on fewer data points. This can be attributed to race estimation being the task in co-training.
Model Data Points White Non-white Average Discrepancy
Age FairFace 100K 60.05 60.63 60.52 0.58
CLIP 400M 62.25 61.95 62.00 -0.30
FaRL 20M 61.49 61.84 61.78 0.35
FaceXFormer 5.9M 58.94 59.44 59.34 0.50
Gender FairFace 100K 94.15 94.41 94.36 0.26
CLIP 400M 94.87 95.78 95.61 0.91
FaRL 20M 95.16 95.77 95.65 0.61
FaceXFormer 5.9M 95.34 95.19 95.22 -0.09
Table 8: Age and gender accuracy w.r.t race groups on FairFace
5.3 Computational Performance Analysis
We present a computational performance analysis of the proposed method compared to previous multi-task models in Table 9 to highlight its efficiency. FaceXFormer achieves the fastest inference speed among multi-task face analysis models, with an FPS of $33.2$ (FP32) and $100.1$ (FP16), outperforming previous multi-task model Faceptor [74]. This improvement is attributed to the proposed FaceX decoder, which employs a novel bi-directional cross-attention mechanism, enabling FaceXFormer to maintain only two decoder layers while ensuring effective face feature extraction. Moreover, FaceXFormer significantly reduces computational cost, requiring only $114$ GFLOPs compared to $167$ GFLOPs in Faceptor, leading to a substantial reduction in latency from $69.9$ ms to $30.1$ ms in FP32 and from $23.7$ ms to $10.0$ ms in FP16. With its reduced computational cost and faster inference, FaceXFormer achieves state-of-the-art performance across most tasks, demonstrating the effectiveness of its lightweight yet powerful design.
Component Model FLOPs Latency (ms) FPS
(GFLOPs) FP32 FP16 FP32 FP16
Backbone Faceptor [74] 64 19.1 6.5 52.3 153.8
FaceXFormer 47 12.5 4.0 80.0 250.0
Decoder Faceptor [74] 85 34.0 11.5 29.4 87.0
FaceXFormer 55 13.6 4.5 73.5 222.2
Task Heads Faceptor [74] 18 6.5 2.2 153.8 454.5
FaceXFormer 12 4.0 1.5 250.0 666.7
Total Inference Time Faceptor [74] 167 69.9 23.7 14.3 42.2
FaceXFormer 114 30.1 10.0 33.2 100.1
Table 9: Computational performance: FaceXFormer vs Faceptor.
6 Conclusion
FaceXFormer introduces a novel end-to-end unified model that efficiently handles a wide range of facial analysis tasks in real-time. By adopting a transformer-based encoder-decoder architecture and representing each task as a learnable token, our approach seamlessly integrates multiple tasks within a single framework while maintaining minimal computational cost and fast inference times. The proposed lightweight decoder, FaceX, incorporates a novel bi-directional cross-attention mechanism, enhancing the model’s ability to learn robust and generalized face representations across diverse tasks. Comprehensive experiments demonstrate that FaceXFormer achieves state-of-the-art performance across multiple facial analysis tasks, achieving a real-time FPS of $33.21$. In broader applications, FaceXFormer can serve as an annotator for large-scale face datasets and can be integrated into existing facial analysis systems to provide extra information, making it a valuable tool for surveillance, subject analysis, and image retrieval.
Acknowledgment
Appendix
Appendix A Overview
  • Broader Impact (Section B)
  • Ablation study (Section C)
  • Cross-dataset Evaluation (Section D)
  • In-the-wild Visualization (Section E)
  • Dataset details (Section F)
Appendix B Discussion
The world is moving towards transformers because of its potential to model large amounts of data [7, 40, 6]. Presently, the face community lacks large-scale annotated datasets to train foundational models capable of performing a wide spectrum of facial tasks. The largest clean dataset, WebFace42M [140], lacks annotations for face parsing, landmarks detection, headpose, expression, race and facial attributes. FaceXFormer can be used as an annotator for large-scale data, and can be continually improved through successive rounds of annotation and fine-tuning. We aim to propel the face community towards developing foundation models that cater to a variety of facial tasks. Additionally, FaceXFormer is a lightweight model that provides real-time output based on task-specific queries and can be appended with existing facial systems to provide additional information. It can also serve as a valuable tool in surveillance, and provide auxiliary information for subject analysis and image retrieval.
Appendix C Ablation Study
To evaluate the impact of different backbones on performance and FPS, we conduct an ablation study comparing various backbone architectures in FaceXFormer. We categorize head pose estimation, landmark prediction, and age estimation as regression (Reg) tasks, while attribute prediction and facial expression recognition fall under classification (Cls). Additionally, face parsing is denoted as segmentation (Seg). The results of these experiments are summarized in Table C.1.
@@ Appendix C Ablation Study
Table C.1: Effect of different backbones on performance and FPS.
@@
From the results, we observe that ConvNeXt achieves the best performance in segmentation with an F1 score of $92.08$%. The Swin Transformer backbone excels in both regression and classification tasks, with a mean error of $4.12$ and a mean accuracy of $90.03$%, respectively. In contrast, MobileNet demonstrates the lowest performance metrics, including an F1 score of $91.21$% and a mean error of $4.64$, highlighting its limitations in handling larger, more complex datasets due to its smaller receptive field compared to the Swin Transformer. The selection of the Swin Transformer as the backbone for FaceXFormer is driven by its superior scalability and global contextual understanding, both of which are essential for facial analysis tasks.
Appendix D Cross-Dataset Evaluation
@@
We conduct additional cross-dataset experiments to demonstrate the effectiveness of FaceXFormer in scenarios that closely resemble real-life conditions. These scenarios involve previously unseen, unconstrained face images characterized by significant variability in background, lighting, pose, and other factors. As shown in Table D.1, FaceXFormer outperforms the existing state-of-the-art model, STARLoss [135], on the 300VW dataset. This highlights FaceXFormer’s effectiveness in landmark detection under in-the-wild scenarios. The cross-dataset results support the rationale presented in this paper: the necessity of a unified facial analysis model capable of performing multiple tasks on unconstrained, in-the-wild faces, particularly for real-time applications. FaceXFormer addresses this gap and achieves state-of-the-art performance.
@@
Table D.1: Cross Dataset evaluation of FaceXFormer.
Appendix E In-the-wild Visualization
We randomly selected images from the web and treated them as ”in-the-wild” images. The qualitative results for all tasks are presented in Figure E.1. Our observations indicate that FaceXFormer produces promising results even in the presence of occlusions, extreme angles, and accessories.
Figure E.1: Visualization of “in-the-wild” images for multiple tasks. Attributes represent the $40$ binary attributes defined in the CelebA [56] dataset, indicating the presence ($1$) or absence (00) of specific facial attributes.
Appendix F Datasets and Implementation Details